I am taking college algebra online and am having problems with equations and inequalities: for example: find all real solutions by factoring:
x^3+4x^2-x-4=0
thanks
Sal

Answer by kiru_khandelwal(79)

x^3 + 4x^2 -x -4 =0
=> x^2(x+4) -(x+4) = 0
=> (x+4)(x^2-1) = 0
=> According to the zero product rule
either x+4 = 0 or x^2-1 = 0
=> if x+4 =0
then x = -4
and if x^2 -1 = 0
then x^2 = 1
=> x = sqrt(1) = +/-1
so x = -4, +/-1